OpenStudy (anonymous):

Using the slope and the y-intercept, graph the line represented by the following equation. 2x - y + 4 = 0

OpenStudy (cwrw238):

first rearrange the equation to slope/intercept form y = mx + c can u do that?

OpenStudy (anonymous):

How would I do that?

OpenStudy (cwrw238):

ok subtract 2x from both sides o the equation 2x - y + 4 = 0 -y + 4 = -2x now subtract 4 from both sides -y = -2x - 4 we need y = mx + c so multiply both sides by -1: y = 2x + 4

OpenStudy (anonymous):

c is where y starts when x is 0. I learned it this way: \[y = mx + y_0 \] m is the slope, since it's positive, the line goes up diagonally from 2 on 2

OpenStudy (cwrw238):

so slope is 2 and intercept on y axis is 4 line will look like 

OpenStudy (anonymous):

Oh wow! Now if the point is on the 4 but it going up...where does the 2 end up??

OpenStudy (cwrw238):

slope is 2 so the line will also go through x = -2